# Logarithmic Enriques varieties

### v1

@@ Abstract.
We introduce logarithmic Enriques varieties as a singular analogue of Enriques manifolds, generalizing the notion of log-Enriques surfaces introduced by Zhang. We focus then on the properties of the subfamily of log-Enriques varieties that admit a quasi-étale cover by a singular symplectic variety and we give many examples.
@@
holomorphic symplectic manifolds, Calabi-Yau varieties, abelian varieties, automorphisms, Enriques varieties
@@ 1. Introduction
In the Enriques–Kodaira classification of smooth complex algebraic surfaces, an Enriques surface is by definition a minimal surface $X$ of Kodaira dimension $\kappa=0$, arithmetic genus $p_{g}(X)\coloneqq h^{0}(K_{X})=0$ and irregularity $q(X)\coloneqq h^{1}(\mathcal{O}_{X})=0$. It follows that the canonical divisor $K_{X}$ is $2$-torsion and that the étale double covering defined by the $2$-torsion line bundle $\mathcal{O}_{X}(K_{X})$ realizes $X$ as the quotient of a K3 surface $Y$ by a fixed point free nonsymplectic involution. As a consequence, its fundamental group $\pi_{1}(X)$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$. The existence of a nonsymplectic automorphism imposes that $Y$ is projective, hence $X$ too. Conversely, every fixed point free involution on a K3 surface $Y$ is nonsymplectic by the holomorphic Lefschetz fixed point formula. This imposes that $Y$ is projective, and the quotient variety is an Enriques surface.
@@
In a different flavour, starting from a smooth complex algebraic surface $X$ with nontrivial but $2$-torsion canonical bundle, the étale degree two covering $Y\to X$ defined by the canonical sheaf is such that $K_{Y}=0$, so $p_{g}(Y)=1$ and $Y$ is a minimal surface which is either a K3 surface or an Abelian surface. If we further assume that the $q(X)=0$, then the Euler characteristic of $X$ is $\chi(X,\mathcal{O}_{X})=1$, so $\chi(Y,\mathcal{O}_{Y})=2$, hence $q(Y)=0$ and $Y$ is a K3 surface. We refer to Beauville [8, Chapter VIII], Barth–Hulek–Peters–van de Ven [6] or Cossec–Dolgachev [23] for more characterizations of Enriques surfaces.
@@
These equivalent points of view on Enriques surfaces offer different strategies to extend their definition in higher dimension in the nonsingular setup, that have been explored in Boissière–Nieper-Wisskirchen–Sarti [15], Oguiso–Schröer [47] and Yoshikawa [54]. In this paper, after reviewing the theory of singular symplectic varieties and of singular Calabi–Yau varieties in Section 2 and after discussing some properties of $K$-torsion varieties in Section 3, we first propose in Definition 4.1 a general definition of Enriques manifolds that interpolates with the previous ones cited above and then we propose in Definition 4.6 an extension of this notion in the singular setup, that we call logarithmic Enriques varieties, following the term logarithmic Enriques surfaces introduced by Zhang [56]. In Section 5 then we discuss general properties and characterizations of these varieties in the special case of log-Enriques varieties of symplectic type, i.e.those log-Enriques varieties admitting a cyclic quasi-étale cover by a symplectic variety. In particular, in Proposition 5.9 we show that if the quasi-étale cover $Y$ is an ISV then the log-Enriques variety $X$ is the quotient of $Y$ by the action of a purely nonsymplectic automorphism. Finally, in Section 6 we give many examples: in particular, in the case of Enriques manifolds, we review the examples already mentioned in the cited references and also construct some new ones.
@@
We have been informed by Angel David Ríos Ortiz, Francesco Denisi, Nikolaos Tsakanikas and Zhixin Xie that they are independently working on an variant of this notion that they called primitive Enriques varieties, that is compatible with our definitions, although the two papers explore different developments.
@@
The authors warmly thank Francesco Denisi, Christian Lehn, Martina Monti, Arvid Perego, Gianluca Piacenza and Giulia Saccà for helpful discussions. The first and the third author have been partially funded by the ANR/DFG projet “Positivity of K-trivial varieties”. The second author received partial support from grants Prin Project 2020 ”Curves, Ricci flat Varieties and their Interactions” and Prin Project 2022 ”Symplectic varieties: their interplay with Fano manifolds and derived categories”; she is a member of the Indam group GNSAGA and was supported by the International Research Laboratory (IRL) LYSM ”Ypatia Laboratory for Mathematical Sciences”.
@@ 2. Preliminaries
it is a divisorial sheaf whose associated linear equivalence class of Weil divisors is the canonical divisor $K_{X}$ of $X$. Following the definition given in [51, p.282] for the divisorial sheaf associated to a Weil divisor, we have $\omega_{X}=\mathcal{O}_{X}(K_{X})$. With a slight, but customary abuse of notation, we denote:
@@
Recall that a normal variety $X$ has rational singularities if for any resolution of singularities $f\colon Z\to X$, one has $R^{>0}f_{\ast}\mathcal{O}_{Z}=0$. By the extension theorem of Kebekus–Schnell [38, Corollary 1.8], if $X$ is a normal variety with rational singularities, the pullback over $X_{\rm reg}$ of any reflexive form extends regularly on any resolution of singularities. We also recall that by the Elkik–Flenner theorem, rational Gorenstein singularities are exactly canonical singularities of index one [52, §3(C) & p.363].
@@
A finite covering map is a finite and surjective morphism $f\colon\widetilde{X}\to X$ between normal projective varieties of the same dimension. It is called quasi-étale if it is étale in codimension one  [30, Definition 3.2&3.3]: this means that there exists a closed subset $Z\subset\widetilde{X}$ of codimension at least two such that the restriction $f\colon\widetilde{X}\setminus Z\to X$ is étale. The augmented irregularity [31, Definition 3.1] of a normal projective variety $X$ is the supremum $\tilde{q}(X)\in\mathbb{N}\cup\{\infty\}$ of the irregularities $q(\widetilde{X})\coloneqq h^{1}(\widetilde{X},\mathcal{O}_{\widetilde{X}})$ for all finite quasi-étale coverings $\widetilde{X}\to X$.
A Calabi–Yau variety (CY) is a normal projective variety $X$ such that $\omega_{X}\cong\mathcal{O}_{X}$, with rational singularities and such that for any finite quasi-étale covering $f\colon\widetilde{X}\to X$, one has $H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{[q]})=0$ for all $0<q<\dim X$ (see [31, Definition 8.16.1]). By Hodge symmetry $\tilde{q}(X)=0$ and $X$ is simply connected if $X$ has even dimension [28, Corollary 13.3]. Nonsingular CY varieties correspond to the CY manifolds as defined in [7].
An irreducible symplectic variety (ISV) is a projective symplectic variety $X$ admitting a symplectic form $\sigma_{X}\in H^{0}(X,\Omega^{[2]}_{X})$ such that that for any finite quasi-étale covering $f\colon\widetilde{X}\to X$, the reflexive pullback $f^{[\ast]}(\sigma_{X})$, as defined in [29, §II.4], generates the exterior algebra of global sections of $\Omega^{[\ast]}_{\widetilde{X}}$ (see [31, Definition 8.16.2] and [3, Definition 1.1]). In particular, we have $H^{0}(X,\Omega^{[2]}_{X})=\mathbb{C}\sigma_{X}$. Clearly $X$ is even-dimensional, $\tilde{q}(X)=0$ and $X$ is simply connected by [28, Corollary 13.3], but $X_{\rm reg}$ is not always simply connected. Nonsingular ISV coincide with the irreducible holomorphic symplectic (IHS) manifolds introduced by Beauville [7].
A primitive symplectic variety (PSV), considered in particular by Schwald [53] and Bakker–Lehn [5], is a projective symplectic variety $X$ such that $q(X)=0$ and $h^{0}(X_{\rm reg},\Omega^{2}_{X})=1$. ISV are PSV but the augmented irregularity of a PSV may be nonzero: take for instance the quotient of an abelian surface by ±$\pm 1$, it is a PSV since its minimal resolution is a Kummer surface but by construction its augmented irregularity is positive. We refer to [4, 49] for a more detailed discussion on these definitions and its possible variants.
@@
The argument for ISV is given in [16, Proposition 3.15], the CY case is similar, we write it for completeness. By the Zariski–Nagata purity theorem [45, 55] (see also [57, Theorem 2.4]), every quasi-étale cover is étale over the nonsingular locus. Since $X_{\rm reg}$ is simply connected, every quasi-étale cover of $X$ is an isomorphism, so we only need to show that $H^{0}(X,\Omega_{X}^{[q]})=0$ for all $0<q<\dim X$. Consider the following diagram:
@@
where $X^{\prime}$ is a resolution of singularities of $X$ and $Z^{\prime}$ is a nonsingular resolution of indeterminacies of the rational dominant map $Z\dashrightarrow X^{\prime}$. For all $0<q<\dim X$, by [39, Proposition 5.8] applied to the dominant morphism $\gamma\colon Z^{\prime}\rightarrow X$, the reflexive pullback morphism $\gamma^{[\ast]}\colon H^{0}(X,\Omega_{X}^{[q]})\to H^{0}(Z^{\prime},\Omega_{Z^{\prime}}^{q})$ is injective. By Hodge symmetry, we have $h^{0}(Z^{\prime},\Omega_{Z^{\prime}}^{q})=h^{q}(Z^{\prime},\mathcal{O}_{Z}^{\prime})$. Since the singularities of $Z$ are rational and since $Z^{\prime}\to Z$ is birational, we have $h^{q}(Z^{\prime},\mathcal{O}_{Z^{\prime}})=h^{q}(Z,\mathcal{O}_{Z})$. By Hodge symmetry [31, Proposition 6.9], since $Z$ has canonical singularities we have $h^{q}(Z,\mathcal{O}_{Z})=h^{0}(Z,\Omega_{Z}^{[q]})$, which is zero since $Z$ is CY. We get $h^{0}(X,\Omega_{X}^{[q]})=0$. ∎
@@ 3. K-torsion varieties
Let $X$ be a normal projective variety. If $X$ has only canonical singularities and if its canonical divisor is numerically trivial, by Kawamata [37, Theorem 8.2] we get that $K_{X}$ is torsion and that $\kappa(X)=0$. Assume now more generally that the singularities of $X$ are log terminal (klt). If $K_{X}$ is numerically effective, the abundance conjecture predicts that some multiple $dK_{X}$ of $K_{X}$ is base-point-free. If we further assume that $\kappa(X)=0$, the morphism induced by the linear system $|dK_{X}|$ maps to a point, so $K_{X}$ is torsion. This motivates the following definition:
@@
Under this assumption, the divisorial sheaf $\omega_{X}$ satisfies $\omega_{X}^{[d]}\cong\mathcal{O}_{X}$, it is thus a reflexive $d$-th root of the structural sheaf. We refer to [40, Definition 2.34] for the definition of klt  singularities. Here we could simply equivalently require log terminal singularities, but the klt  framework is more common.
Following [40, Definition 5.19], if $X$ is a $K$-torsion variety of index $d$, the quasi-étale cyclic covering associated to $\omega_{X}$, with group $\mu_{d}$ (the order $d$ cyclic group), is:
@@
Since $\omega_{X}$ is locally free on $X_{\rm reg}$, the restriction of $p$ to $p^{-1}(X_{\rm reg})\to X_{\rm reg}$ is étale, with $p^{-1}(X_{\rm reg})\subset Y_{\rm reg}$. Following Kawamata [37, Section 8], we call $Y$ the canonical cover of $X$. Recall that $Y$ is projective, irreducible, normal (this is a consequence of [51, Proposition 2]), it has canonical singularities [40, Corollary 5.21] and $K_{Y}=0$. More generally, if $X$ is klt  and $p\colon Y\to X$ is any quasi-étale morphism with $Y$ normal, then $Y$ is also klt  by [40, Proposition 5.20].
@@
Let $p\colon Y\to X$ be a quasi-étale cyclic covering of degree $d\geq 2$ between projective normal varieties. We have the following exact sequence of class groups:
@@
This statement extends [47, Proposition 2.8] to the singular setup, using similar arguments. By the Zariski–Nagata purity theorem, since $p$ is quasi-étale, it is étale over the nonsingular locus $X_{\rm reg}$ and $Y^{\circ}\coloneqq p^{-1}(X_{\rm reg})$ is an open subset of $Y_{\rm reg}$ whose complementary in $Y$ has codimension at least two. The restriction $p\colon Y^{\circ}\to X_{\rm reg}$ is an étale cyclic cover with group $\mu_{d}$. Following the methods and notation in [33, §5.2], there are two spectral sequences associated to $p$, applied here to the sheaf $\mathcal{O}_{Y^{\circ}}^{\ast}$:
@@
Proof of assertion (i). The subgroup $\mathrm{Cl}_{a}(Y)$ of $\mathrm{Cl}(Y)$ parametrizing linear equivalence classes of Weil divisors on $Y$ algebraically equivalent to zero admits a structure of abelian variety defined by Lang [41, IV§4] as the Picard variety $\mathrm{P}(Y)\coloneqq\mathrm{Alb}(Y)^{\vee}$, the dual of the Albanese variety of $Y$ (see for instance [13] for more details). The dimension of the Picard variety is thus $q(Y)=h^{1}(Y,\mathcal{O}_{Y})=0$ by assumption, so $\mathrm{Cl}_{a}(Y)=\{0\}$. Since the Néron–Severi group $\mathrm{NS}(Y)\coloneqq\mathrm{Cl}(Y)/\mathrm{Cl}_{a}(Y)$ is finitely generated by [36, Théorème 3], we obtain that $\mathrm{Cl}(Y)$ is finitely generated. Using the exact sequence (1) we get that $\mathrm{Cl}(X)$ is also finitely generated.
Proof of assertion (ii). To prove that $dK_{X}=0$, it is enough to show that $dK_{X_{\rm reg}}=0$ since $X$ is normal, so after restricting to the étale covering $p\colon Y^{\circ}\to X_{\rm reg}$, we may assume that $Y$ and $X$ are nonsingular. Since $p$ is étale, we have $p^{\ast}\mathcal{O}_{X}(K_{X})=\mathcal{O}_{Y}(K_{Y})=\mathcal{O}_{Y}$ and the result follows using the exact sequence (1). Here is a more direct argument that will be used below. Since $p$ is finite and $X$ is nonsingular, $p$ is flat so $p_{\ast}\mathcal{O}_{Y}$ is locally free of rank $d$. By projection formula we get:

### v2

@@ Abstract.
We introduce logarithmic Enriques varieties as a singular analogue of Enriques manifolds, generalizing the notion of log-Enriques surfaces introduced by Zhang. We focus mainly on the properties of the subfamily of log-Enriques varieties that admit a quasi-étale cover by a singular symplectic variety and we give many examples.
@@
holomorphic symplectic manifolds, symplectic singularities, Calabi-Yau varieties, abelian varieties, automorphisms, Enriques varieties
@@ 1. Introduction
In the Enriques–Kodaira classification of smooth complex algebraic surfaces, an Enriques surface is by definition a minimal surface $X$ of Kodaira dimension $\kappa=0$, arithmetic genus $p_{g}(X)\coloneqq h^{0}(K_{X})=0$ and irregularity $q(X)\coloneqq h^{1}(\mathcal{O}_{X})=0$. It follows that the canonical divisor $K_{X}$ is $2$-torsion and that the étale double cover defined by the $2$-torsion line bundle $\mathcal{O}_{X}(K_{X})$ realizes $X$ as the quotient of a K3 surface $Y$ by a fixed point free nonsymplectic involution. As a consequence, its fundamental group $\pi_{1}(X)$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$. The existence of a nonsymplectic automorphism imposes that $Y$ is projective, hence $X$ too. Conversely, every fixed point free involution on a K3 surface $Y$ is nonsymplectic by the holomorphic Lefschetz fixed point formula. This imposes that $Y$ is projective, and the quotient variety is an Enriques surface.
@@
In a different flavour, starting from a smooth complex algebraic surface $X$ with nontrivial but $2$-torsion canonical bundle, the étale degree two cover $Y\to X$ defined by the canonical sheaf is such that $K_{Y}=0$, so $p_{g}(Y)=1$ and $Y$ is a minimal surface which is either a K3 surface or an Abelian surface. If we further assume that $q(X)=0$, then the Euler characteristic of $X$ is $\chi(X,\mathcal{O}_{X})=1$, so $\chi(Y,\mathcal{O}_{Y})=2$, hence $q(Y)=0$ and $Y$ is a K3 surface. We refer to Beauville [8, Chapter VIII], Barth–Hulek–Peters–van de Ven [6] or Cossec–Dolgachev [25] for more characterizations of Enriques surfaces.
@@
These equivalent points of view on Enriques surfaces offer different strategies to extend their definition in higher dimension in the nonsingular setup, that have been explored in Boissière–Nieper-Wisskirchen–Sarti [17], Oguiso–Schröer [51] and Yoshikawa [59]: these are called Enriques manifolds. A recent interest for Enriques manifolds occurs in the framework of the investigation of the Morrison–Kawamata cone conjecture by Pacienza and the third author [53]. In light of recent developments on the Beauville–Bogomolov decomposition of singular varieties with trivial first Chern class [29, 35, 39], it is interesting to extend further the definition of Enriques manifolds to the singular setting. In this paper, after reviewing the theory of singular symplectic varieties and of singular Calabi–Yau varieties in Section 2, and after discussing some properties of $K$-torsion varieties in Section 3, we first propose in Definition 4.1 a general definition of Enriques manifolds that interpolates with the previous ones cited above and then we propose in Definition 4.6 an extension of this notion in the singular setup, that we call logarithmic Enriques varieties, following the term logarithmic Enriques surfaces introduced by Zhang [61]. Then in Section 5 we discuss general properties and characterizations of these varieties in the special case of log-Enriques varieties of symplectic type, i.e. those log-Enriques varieties admitting a cyclic quasi-étale cover by a symplectic variety. In particular, in Proposition 5.9 we show that if the quasi-étale cover $Y$ is a PSV (respectively an ISV) then the log-Enriques variety $X$ is the quotient of $Y$ by the action of a purely nonsymplectic automorphism. Finally, in Section 6 we give many examples: in particular, in the case of Enriques manifolds, we review the examples already mentioned in the cited references and we also construct some new ones.
@@
Angel David Ríos Ortiz, Francesco Denisi, Nikolaos Tsakanikas and Zhixin Xie [27] have been working independently and simultaneously on a variant of this notion that they called primitive Enriques varieties, that is compatible with our definitions, although they explore different developments as ours.
@@
The authors warmly thank Francesco Denisi, Christian Lehn, Martina Monti, Gianluca Pacienza, Arvid Perego, Giulia Saccà and Nikolaos Tsakanikas for helpful discussions. The first and the third author have been partially funded by the ANR/DFG project ANR-23-CE40-0026 “Positivity on K-trivial varieties”. The second author received partial support from grants Prin Project 2020 ”Curves, Ricci flat Varieties and their Interactions” and Prin Project 2022 ”Symplectic varieties: their interplay with Fano manifolds and derived categories”; she is a member of the Indam group GNSAGA and was supported by the International Research Laboratory (IRL) LYSM ”Ypatia Laboratory for Mathematical Sciences”.
@@ 2. Preliminaries
it is a divisorial sheaf whose associated linear equivalence class of Weil divisors is the canonical divisor $K_{X}$ of $X$. Following the definition given in [56, p.282] for the divisorial sheaf associated to a Weil divisor, we have $\omega_{X}=\mathcal{O}_{X}(K_{X})$. With a slight, but customary abuse of notation, we denote:
@@
Recall that a normal variety $X$ has rational singularities if for any resolution of singularities $f\colon Z\to X$, one has $R^{>0}f_{\ast}\mathcal{O}_{Z}=0$. By the extension theorem of Kebekus–Schnell [42, Corollary 1.8], if $X$ is a normal variety with rational singularities, the pullback over $X_{\rm reg}$ of any reflexive form extends regularly on any resolution of singularities. We also recall that by the Elkik–Flenner theorem, rational Gorenstein singularities are exactly canonical singularities of index one [57, §3(C) & p.363].
@@
A finite cover map is a finite and surjective morphism $f\colon\widetilde{X}\to X$ between normal projective varieties of the same dimension. It is called quasi-étale if it is étale in codimension one  [34, Definition 3.2&3.3]: this means that there exists a closed subset $Z\subset\widetilde{X}$ of codimension at least two such that the restriction $f\colon\widetilde{X}\setminus Z\to X$ is étale. The augmented irregularity [35, Definition 3.1] of a normal projective variety $X$ is the supremum $\tilde{q}(X)\in\mathbb{N}\cup\{\infty\}$ of the irregularities $q(\widetilde{X})\coloneqq h^{1}(\widetilde{X},\mathcal{O}_{\widetilde{X}})$ for all finite quasi-étale covers $\widetilde{X}\to X$.
A Calabi–Yau variety (CY) is a normal projective variety $X$ such that $\omega_{X}\cong\mathcal{O}_{X}$, with rational singularities and such that for any finite quasi-étale cover $f\colon\widetilde{X}\to X$, one has $H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{[q]})=0$ for all $0<q<\dim X$ (see [35, Definition 8.16.1] or [32, Definition 1.3]. By Hodge symmetry, $\tilde{q}(X)=0$ and $X$ is simply connected if $X$ has even dimension [32, Corollary 13.3]. Nonsingular CY varieties correspond to the CY manifolds as defined in [7]. A weaker variant co-exists in the literature, without the condition on all étale covers, assuming only that $X$ is normal with rational singularities, $\omega_{X}\cong\mathcal{O}_{X}$ and $H^{0}(X,\Omega_{X}^{[q]})=0$ for all $0<q<\dim X$: to avoid confusion, we call here these later weak CY varieties.
An irreducible symplectic variety (ISV) is a projective symplectic variety $X$ admitting a symplectic form $\sigma_{X}\in H^{0}(X,\Omega^{[2]}_{X})$ such that for any finite quasi-étale cover $f\colon\widetilde{X}\to X$, the reflexive pullback $f^{[\ast]}(\sigma_{X})$, as defined in [33, §II.4], generates the exterior algebra of global sections of $\Omega^{[\ast]}_{\widetilde{X}}$ (see [35, Definition 8.16.2] and [3, Definition 1.1]). In particular, we have $H^{0}(X,\Omega^{[2]}_{X})=\mathbb{C}\sigma_{X}$. Clearly $X$ is even-dimensional, $\tilde{q}(X)=0$ and $X$ is simply connected by [32, Corollary 13.3], but $X_{\rm reg}$ is not always simply connected. Nonsingular ISV coincide with the irreducible holomorphic symplectic (IHS) manifolds introduced by Beauville [7].
A primitive symplectic variety (PSV), as considered in particular by Schwald [58] and Bakker–Lehn [5], is a projective symplectic variety $X$ such that $q(X)=0$ and $H^{0}(X,\Omega^{[2]}_{X})$ is one-dimensional. ISV are PSV but the augmented irregularity of a PSV may be nonzero: take for instance the quotient of an abelian surface by ±$\pm 1$, it is a PSV since its minimal resolution is a Kummer surface but by construction its augmented irregularity is positive. We refer to [4, 54] for a more detailed discussion on these definitions and its possible variants.
@@
The argument for ISV is given in [18, Proposition 3.15], the CY case is similar, we write it for completeness. By the Zariski–Nagata purity theorem [49, 60] (see also [62, Theorem 2.4]), every quasi-étale cover is étale over the nonsingular locus. Since $X_{\rm reg}$ is simply connected, every quasi-étale cover of $X$ is an isomorphism by the Zariski Main Theorem, so we only need to show that $H^{0}(X,\Omega_{X}^{[q]})=0$ for all $0<q<\dim X$. Consider the following diagram:
@@
where $X^{\prime}$ is a resolution of singularities of $X$ and $Z^{\prime}$ is a nonsingular resolution of indeterminacies of the rational dominant map $Z\dashrightarrow X^{\prime}$. For all ${0<q<\dim X}$, by [43, Proposition 5.8] applied to the dominant morphism $\gamma\colon Z^{\prime}\rightarrow X$, the reflexive pullback morphism $\gamma^{[\ast]}\colon H^{0}(X,\Omega_{X}^{[q]})\to H^{0}(Z^{\prime},\Omega_{Z^{\prime}}^{q})$ is injective. By Hodge symmetry [35, Proposition 6.9], we have $h^{0}(Z^{\prime},\Omega_{Z^{\prime}}^{q})=h^{q}(Z^{\prime},\mathcal{O}_{Z}^{\prime})$. Since the singularities of $Z$ are rational and since $Z^{\prime}\to Z$ is birational, we have $h^{q}(Z^{\prime},\mathcal{O}_{Z^{\prime}})=h^{q}(Z,\mathcal{O}_{Z})$. By Hodge symmetry again, since $Z$ has canonical singularities we have $h^{q}(Z,\mathcal{O}_{Z})=h^{0}(Z,\Omega_{Z}^{[q]})$, which is zero since $Z$ is CY. We get $h^{0}(X,\Omega_{X}^{[q]})=0$. ∎
@@ 3. K-torsion varieties
Let $X$ be a normal projective variety. If $X$ has only canonical singularities and if its canonical divisor is numerically trivial, by Kawamata [41, Theorem 8.2] we get that $K_{X}$ is torsion and that $\kappa(X)=0$. Assume now more generally that the singularities of $X$ are log terminal (klt). If $K_{X}$ is numerically effective, the abundance conjecture predicts that some multiple $dK_{X}$ of $K_{X}$ is base-point-free. If we further assume that $\kappa(X)=0$, the morphism induced by the linear system $|dK_{X}|$ maps to a point, so $K_{X}$ is torsion. This motivates the following definition:
@@
Under this assumption, the divisorial sheaf $\omega_{X}$ satisfies $\omega_{X}^{[d]}\cong\mathcal{O}_{X}$, it is thus a reflexive $d$-th root of the structural sheaf. We refer to [44, Definition 2.34] for the definition of klt  singularities. Here we could simply equivalently require log terminal singularities, but the klt  framework is more common.
Following [44, Definition 5.19], if $X$ is a $K$-torsion variety of index $d$, the quasi-étale cyclic cover associated to $\omega_{X}$, with group $\mu_{d}$ (the order $d$ cyclic group), is:
@@
Since $\omega_{X}$ is locally free on $X_{\rm reg}$, the restriction of $p$ to $p^{-1}(X_{\rm reg})\to X_{\rm reg}$ is étale, with $p^{-1}(X_{\rm reg})\subset Y_{\rm reg}$. Following Kawamata [41, Section 8], we call $Y$ the canonical cover of $X$. Recall that $Y$ is projective, irreducible, normal (this is a consequence of [56, Proposition 2]), it has canonical singularities [44, Corollary 5.21] and $K_{Y}=0$. More generally, if $X$ is klt  and $p\colon Y\to X$ is any quasi-étale morphism with $Y$ normal, then $Y$ is also klt  by [44, Proposition 5.20].
@@
Let $p\colon Y\to X$ be a quasi-étale cyclic cover of degree $d\geq 2$ between projective normal varieties. We have the following exact sequence of class groups:
@@
This statement extends [51, Proposition 2.8] to the singular setup, using similar arguments. By the Zariski–Nagata purity theorem, since $p$ is quasi-étale, it is étale over the nonsingular locus $X_{\rm reg}$ and $Y^{\circ}\coloneqq p^{-1}(X_{\rm reg})$ is an open subset of $Y_{\rm reg}$ whose complementary in $Y$ has codimension at least two. The restriction $p\colon Y^{\circ}\to X_{\rm reg}$ is an étale cyclic cover with group $\mu_{d}$. Following the methods and notation in [37, §5.2], there are two spectral sequences associated to $p$, applied here to the sheaf $\mathcal{O}_{Y^{\circ}}^{\ast}$:
@@
Proof of assertion (i). The subgroup $\mathrm{Cl}_{a}(Y)$ of $\mathrm{Cl}(Y)$ parametrizing linear equivalence classes of Weil divisors on $Y$ algebraically equivalent to zero admits a structure of abelian variety defined by Lang [45, IV§4] as the Picard variety $\mathrm{P}(Y)\coloneqq\mathrm{Alb}(Y)^{\vee}$, the dual of the Albanese variety of $Y$ (see for instance [15] for more details). The dimension of the Picard variety is thus $q(Y)=h^{1}(Y,\mathcal{O}_{Y})=0$ by assumption, so $\mathrm{Cl}_{a}(Y)=\{0\}$. Since the Néron–Severi group $\mathrm{NS}(Y)\coloneqq\mathrm{Cl}(Y)/\mathrm{Cl}_{a}(Y)$ is finitely generated by [40, Théorème 3], we obtain that $\mathrm{Cl}(Y)$ is finitely generated. Using the exact sequence (1) we get that $\mathrm{Cl}(X)$ is also finitely generated.
Proof of assertion (ii). To prove that $dK_{X}=0$, it is enough to show that ${dK_{X_{\rm reg}}=0}$ since $X$ is normal, so after restricting to the étale cover ${p\colon Y^{\circ}\to X_{\rm reg}}$, we may assume that $Y$ and $X$ are nonsingular. Since $p$ is étale, we have $p^{\ast}\mathcal{O}_{X}(K_{X})=\mathcal{O}_{Y}(K_{Y})=\mathcal{O}_{Y}$ and the result follows using the exact sequence (1). Here is a more direct argument that will be used below. Since $p$ is finite and $X$ is nonsingular, $p$ is flat so $p_{\ast}\mathcal{O}_{Y}$ is locally free of rank $d$. By projection formula we get: